# On the Compositional Generalization of Multimodal LLMs for Medical Imaging

### v1

@@ Abstract
Multimodal large language models (MLLMs) hold significant potential in the medical field, but their capabilities are often limited by insufficient data in certain medical domains, highlighting the need for understanding what kinds of images can be used by MLLMs for generalization. Current research suggests that multi-task training outperforms single-task as different tasks can benefit each other, but they often overlook the internal relationships within these tasks, providing limited guidance on selecting datasets to enhance specific tasks. To analyze this phenomenon, we attempted to employ compositional generalization (CG)—the ability of models to understand novel combinations by recombining learned elements—as a guiding framework. Since medical images can be precisely defined by Modality, Anatomical area, and Task, naturally providing an environment for exploring CG. Therefore, we assembled 106 medical datasets to create Med-MAT for comprehensive experiments. The experiments confirmed that MLLMs can use CG to understand unseen medical images and identified CG as one of the main drivers of the generalization observed in multi-task training. Additionally, further studies demonstrated that CG effectively supports datasets with limited data and delivers consistent performance across different backbones, highlighting its versatility and broad applicability. Med-MAT is publicly available at github.com/FreedomIntelligence/Med-MAT.
@@
On the Compositional Generalization of Multimodal LLMs
@@ 1 Introduction
Multimodal large language models (MLLMs) Liu et al. (2023); Li et al. (2024); Chen et al. (2024b) are showing great promise for the medical community, facilitating efficient consultations for doctors and providing patients with anytime access to their medical conditions. However, limited data on rare or privacy-restricted diseases often restricts MLLMs’ performance, making it crucial to explore what kinds of images can be used by MLLMs to generalize these diseases or learn them efficiently.
Current research on medical image generalization Mo and Liang (2024); Ren et al. (2024) has demonstrated that models trained on multiple tasks outperform those trained on a single task as they can leverage potential knowledge from other tasks. However, these studies have not conducted a detailed analysis of which data within multitask datasets can mutually complement each other. Consequently, we aim to investigate the phenomenon of mutual improvement in MLLMs’ understanding of medical images from the perspective of composition generalization.
@@
Compositional generalization (CG) Xu et al. (2022); Li et al. (2019) refers to a model’s ability to learn fundamental elements and recombine them in novel ways to understand unseen combinations (Figure 1). In medical imaging, each image can be categorized by Modality, Anatomical area, and medical Task, presenting numerous natural opportunities for CG. We defined these three elements as the MAT-Triplet and collected 106 medical datasets, subsequently merging those that share the same MAT-Triplet to create the Med-MAT dataset. Ultimately, Med-MAT comprises 53 subsets, encompassing 11 modalities, 14 anatomical regions, and 13 medical tasks, providing a foundation for investigating CG and other generalization methods.
@@
To verify the existence of CG, we designated certain datasets as Target data and selected all the Related data from Med-MAT that shared the same MAT-Triplet with the Target data for generalization. In addition, we conducted larger-scale experiments and observed the changes in model generalization performance after deliberately disrupting CG. Ultimately, the experiments demonstrated that MLLMs can leverage CG to understand unseen medical images and CG is one of the key factors driving the generalization observed in multi-task training.
Subsequently, the practical applications of CG and its effects on different backbones were further explored. We found that increasing the volume of CG combinations continuously enhanced the model’s understanding of the Target data. While CG’s direct impact on unseen data was not always apparent, CG could assist the model in efficient fitting with limited data by being incorporated into the training data of the Target. Furthermore, we discovered that CG exists in different MLLM backbones, underscoring CG’s broad applicability.
@@
Here are the key contributions of our work: 1) A VQA dataset, Med-MAT, has been constructed, providing a platform to explore the generalization of MLLMs on medical images. 2) Through this dataset, we observed that MLLMs can utilize compositional generalization to understand unseen images, demonstrating that this is one of the main forms of generalization for medical MLLMs. 3) Finally, the contribution of CG to understanding medical domains with limited data has been further explored, with its presence across different backbones demonstrating its broad applicability.
| Model | 02 | 03 | 07 | 08 | 09 | 11 | 13 | 14 | 15 | 16 | 18 | 19 | 21 | 22 | 23 | 25 | 26 | 28 | 30 | 31 | 32 | 33 | 35 | 36 | 37 |
| Baseline | 22 | 47 | 40 | 25 | 26 | 27 | 28 | 24 | 22 | 24 | 25 | 23 | 49 | 26 | 25 | 24 | 49 | 30 | 49 | 21 | 49 | 20 | 25 | 23 | 19 |
@@ 2 Med-MAT
Most existing datasets for MLLMs Zhang et al. (2023c); Li et al. (2024); Chen et al. (2024b), primarily VQA datasets, provide broad coverage but lack attribute annotations for individual samples, which are not suitable for CG exploration. To address this gap, we curated a large collection of image-text pairs to develop Med-MAT, ensuring that each sample is explicitly defined by MAT-Triplet. Section 2.1 provides data processing methods; and Section 2.2 presents a pilot experiment.
2.1 Data Processing
@@
Data Construction: Med-MAT contains a total of 106 image-label pair medical datasets, sourced from various medical public challenges or high-quality annotated datasets. All datasets are categorized according to their MAT-Triplet, with data having identical elements grouped into a single subset (Figure 2). Labels are manually clustered to ensure that annotations with the same meaning are not repeatedly used. In total, Med-MAT covers 11 medical modalities, 14 anatomical areas, and 13 medical tasks, hoping that it can spread across various medical tasks like a mat. (Data lists are shown in Appendix B)
@@
Data Distribution: All subsets are divided into training and test sets following their original distributions or using a 9:1 ratio. To ensure a fair comparison, each training set is limited to 3,000 samples, with label balance maintained as much as possible. Any subset that cannot meet this requirement is treated as an OOD (out-of-distribution) dataset. For the test sets, we strictly balance the number of samples per label to ensure that the accuracy metric reliably reflects model performance.
@@
QA Pairs Construction: To enable MLLMs to directly train and test on Med-MAT, all image-label paired data were converted into a visual question-answering (VQA) format (Figure 3). Specifically, each subset was manually assigned 6 instructions to guide the MLLM in answering the subset task. For convenience, all samples were converted into single-choice questions with up to four options, and the remaining distractor options were randomly drawn from other labels within the subset. To mitigate potential evaluation biases arising from varying option counts, the ImageWikiQA dataset Zhang et al. (2024b), a non-medical dataset consisting of single-answer, four-option questions, was incorporated during the training.
2.2 A Pilot Study of Data Composition
@@
Experiments Setup: Experiments in this paper mainly focused on classification datasets to explore the model’s combinatorial generalization capabilities across various image types. The base model LLaVA-v1.5-7B-Vicuna Liu et al. (2023) features a transparent pretraining process and uses very little medical data during training, thereby minimizing the risk of knowledge leakage. By leveraging the flexibility of MLLM, we achieved task switching and generalization simply by modifying the corresponding prompts, significantly simplifying generalization studies. Each experiment was conducted over 5 epochs using 8 A800 (80GB) GPUs, with a batch size of 32 and a learning rate of 5e-6. Generalization performance was assessed by comparing the model’s accuracy on the Target data.
Pilot Study: To explore the benefits of data composition for downstream tasks, all in-distribution (ID) datasets were combined for multi-task training, with single-task training on individual ID datasets serving as the control. The results (Table 1 and 2) showed that multi-task training outperformed single-task training on specific tasks, and also surpassed the baseline in predicting unseen OOD datasets. This suggests that certain data combinations can enhance the model’s classification performance, highlighting the potential for CG. Exploring which combinations provide valuable insights for downstream medical tasks in a promising avenue for further research.
Take-away 1: Certain data combinations can assist MLLMs in medical classification tasks.
@@ 3 Proof of Concept on CG
To explore whether MLLMs can leverage CG to understand unseen images, controlled variable studies were first conducted using all possible CG combinations to determine if CG exists among the MAT-Triplet (Section 3.1). Subsequently, the number of combinations was increased to explore more generalizable insights, aiming to assess the performance gains of multi-task training (Section 3.2).
| Related Combination | Target Subset | Baseline | Baseline+ | Trained | |
| Lung, COVID | Brain, Cancer | Lung, Cancer | 25 | 25 | 27 |
| Lung, Cancer | Brain, State | Lung, State | 47 | 46 | 50 |
| Brain, Cancer | Lung, State | Brain, State | 33 | 50 | 57 |
| Bones, Level | Lung, State | Bones, State | 49 | 53 | 51 |
| Bones, Level | Brain, State | Bones, State | 49 | 53 | 72 |
| Bones, Level | Breast, Diseases | Bones, Diseases | 37 | 33 | 39 |
| Bones, Level | Lung, Diseases | Bones, Diseases | 37 | 33 | 43 |
| Bones, Level | Chest, Diseases | Bones, Diseases | 37 | 31 | 43 |
| Bones, State | Breast, Diseases | Bones, Diseases | 37 | 37 | 43 |
| Bones, State | Lung, Diseases | Bones, Diseases | 37 | 37 | 43 |
| Bones, State | Chest, Diseases | Bones, Diseases | 37 | 37 | 41 |
| Lung, COVID | Breast, Diseases | Lung, Diseases | 49 | 48 | 51 |
| Lung, COVID | Bones, Diseases | Lung, Diseases | 49 | 48 | 52 |
| Lung, COVID | Chest, Diseases | Lung, Diseases | 49 | 48 | 51 |
| CT, Cancer | X-ray, COVID | CT, COVID | 47 | 46 | 72 |
| CT, COVID | X-ray, Diseases | X-ray, COVID | 30 | 21 | 49 |
| CT, State | X-ray, Diseases | X-ray, State | 30 | 21 | 46 |
| CT, State | X-ray, Cancer | CT, Cancer | 33 | 28 | 28 |
| CT, Brain(State) | X-ray, Bones | X-ray, Brain | 49 | 49 | 91 |
| CT, Brain | X-ray, Lung | X-ray, Brain | 49 | 50 | 81 |
| CT, Brain(Cancer) | X-ray, Bones | X-ray, Brain | 25 | 51 | 74 |
| CT, Brain | X-ray, Lung | X-ray, Brain | 49 | 52 | 52 |
| X-ray, Brain | CT, Lung(State) | CT, Brain(State) | 33 | 50 | 60 |
| X-ray, Lung | CT, Brain | CT, Lung(Cancer) | 25 | 25 | 36 |
| X-ray, Lung | CT, Brain(State) | CT, Lung | 47 | 50 | 81 |
| X-ray, Lung | CT, Brain(Cancer) | CT, Lung | 47 | 50 | 71 |
| CT, Lung (State) | X-ray, Bones | X-ray, Lung | 30 | 32 | 28 |
| CT, Lung (State) | X-ray, Brain | X-ray, Lung | 30 | 32 | 35 |
| CT, Lung (Cancer) | X-ray, Bones | X-ray, Lung | 30 | 32 | 41 |
| CT, Lung (Cancer) | X-ray, Brain | X-ray, Lung | 30 | 32 | 42 |
| Der, Skin, Cancer | FP, Fundus, Diseases | Der, Skin, Diseases | 25 | 29 | 33 |
| Der, Skin, Cancer | OCT, Retine, Diseases | Der, Skin, Diseases | 25 | 29 | 33 |
| Der, Skin, Diseases | DP, Mouth, Cancer | Der, Skin, Cancer | 40 | 33 | 63 |
| Der, Skin, Diseases | Mic, Cell, Cancer | Der, Skin, Cancer | 40 | 33 | 63 |
| DP, Mouth, State | Der, Skin, Cancer | DP, Mouth, Cancer | 48 | 50 | 52 |
| DP, Mouth, State | Mic, Cell, Cancer | DP, Mouth, Cancer | 48 | 50 | 55 |
| FP, Fundus, Diseases | Mic, Cell, Level | FP, Fundus, Level | 33 | 36 | 42 |
| Mic, Cell, Cell Identification | FP, Fundus, Level | Mic, Cell, Level | 23 | 33 | 32 |
| Mic, Cell, Cell identification | Der, Skin, Cancer | Mic, Cell, Cancer | 49 | 50 | 50 |
| Mic, Cell, Cell identification | DP, Mouth, Cancer | Mic, Cell, Cancer | 49 | 51 | 62 |
| Mic, Cell, Level | Der, Skin, Cancer | Mic, Cell, Cancer | 49 | 51 | 52 |
| Mic, Cell, Level | DP, Mouth, Cancer | Mic, Cell, Cancer | 49 | 51 | 58 |
| Mic, Cell, Cancer | FP, Fundus, Level | Mic, Cell, Level | 23 | 24 | 27 |
3.1 Controlled Variable Study on CG
3.1.1 Experiment Setup
@@
In this section, the existence of CG was explored from a finer perspective, focusing on CG with only two elements varying while the third remained constant. For example, we fixed the Modality and examined the CG between Areas and Tasks. Additionally, we identified specific Modality-Area pairs, such as dermoscopy paired consistently with skin, which were treated as a special category. These 4 different fixed formats were classified into distinct Direction Types.
@@
Following the setup in Section 2.2, we evaluated the model’s performance on Target data. Baseline refers to the model without any training, while Trained refers to the model trained solely on Related data. To ensure that our conclusions are not influenced by the amount of training data, we randomly sampled an equal number of data from the Unrelated subsets, and this configuration is referred to as Baseline+.
3.1.2 Analysis of CG in MLLMs
@@
Results are shown in Table 3 and it can be observed that almost all CG combinations are able to generalize to downstream tasks, highlighting that MLLMs can leverage CG to generalize Target data across all Direction Types. Besides that, since this experiment focused solely on two-element tuples, we further investigated three-element tuples in Section 4.3, where we also observed similarly strong generalizations when obtaining MAT-Triplet elements from three different datasets.
@@
Take-away 2: MLLMs can leverage CG to understand unseen medical images.
@@
Take-away 3: Generalization only arises from MAT-Triplet related medical datasets.
3.1.3 Analysis of Non-obvious Cases
Some Trained models show minimal gains or performance declines, especially in Level or Diseases Identification tasks. The images in the Level Identification are very similar, while the Diseases Identification involves a large number of target labels, both of which require large datasets for training, which may explain the poor performance. To further analyze them, we modified the evaluation scenario to use CG combinations for assisting in Target data training with limited samples (Section 4.2).
3.2 Scaling the Combination Number of CG
To ensure the universality of CG, we expanded the number of combination datasets to evaluate its effects on a larger scale. This section explores the following questions: (Q1) While Meta CG experiments indicate that Unrelated combinations provide no benefit to Target data, can generalization arise when training incorporates more Unrelated combinations, simulating a multi-task scenario? (Q2) Previous studies suggest that multi-task training generally promotes better generalization than single-task training. To what extent does this generalization originate from CG?
3.2.1 Experiment Setup
Selection Strategy: To ensure a balanced evaluation of Related and Unrelated combinations, Subset 03 and Subset 28 were chosen as Target datasets because they exhibit the most balanced ratios of Related to Unrelated subsets (13:11 for Subset 03 and 11:13 for Subset 28), making them ideal for providing a diverse range of compositions in the scale-up experiments.

### v2

# Exploring Compositional Generalization of Multimodal LLMs for Medical Imaging

@@ Abstract
Medical imaging provides essential visual insights for diagnosis, and multimodal large language models (MLLMs) are increasingly utilized for its analysis due to their strong generalization capabilities; however, the underlying factors driving this generalization remain unclear. Current research suggests that multi-task training outperforms single-task as different tasks can benefit each other, but they often overlook the internal relationships within these tasks. To analyze this phenomenon, we attempted to employ compositional generalization (CG), which refers to the models’ ability to understand novel combinations by recombining learned elements, as a guiding framework. Since medical images can be precisely defined by Modality, Anatomical area, and Task, naturally providing an environment for exploring CG, we assembled 106 medical datasets to create Med-MAT for comprehensive experiments. The experiments confirmed that MLLMs can use CG to understand unseen medical images and identified CG as one of the main drivers of the generalization observed in multi-task training. Additionally, further studies demonstrated that CG effectively supports datasets with limited data and confirmed that MLLMs can achieve CG across classification and detection tasks, underscoring its broader generalization potential. Med-MAT is available at https://
github.com/FreedomIntelligence/Med-MAT.
@@
Exploring Compositional Generalization of Multimodal LLMs
@@ 1 Introduction
Medical imaging provides essential visual insights into the structures of the human body, making it a critical tool for medical diagnosis. Recently, multimodal large language models (MLLMs) Liu et al. (2023); Li et al. (2024); Chen et al. (2024b) have been employed to analyze these images due to their strong interpretability and generalization capabilities. In this paper, we focus on the latter: generalization of MLLMs in medical imaging. Current research Mo and Liang (2024); Ren et al. (2024) has demonstrated that models trained on multiple tasks outperform those trained on a single task as they can leverage potential knowledge from other tasks. Yet, the underlying factors that contribute to this generalization remain insufficiently explored.
To this end, we take the perspective of composition generalization (CG) Li et al. (2019); Xu et al. (2022); Tang et al. (2024) to investigate the generalization phenomenon of mutual improvement in MLLMs’ understanding of medical images. Specifically, CG is the model’s ability to learn fundamental elements and recombine them in novel ways to understand unseen combinations (e.g., learning Cat from White Cat and Black from Black Dog, then generalizing to Black Cat, as shown in Figure 1).
In this paper, we categorize each image to three elements: Modality
, Anatomical area
, and medical Task
@@
, presenting numerous natural opportunities for CG. We defined these three elements as the MAT-Triplet and collected 106 medical datasets, subsequently merging those that share the same MAT-Triplet to create the Med-MAT dataset. Ultimately, Med-MAT comprises 53 subsets, encompassing 11 modalities, 14 anatomical regions, and 13 medical tasks, providing a foundation for investigating CG and other generalization methods.
@@
To verify the existence of CG, we designated specific datasets as Target data and selected all Related data from Med-MAT that shared the same MAT-Triplet with the Target data. Using these data combinations, we accessed the generalization performance of MLLMs and observed that they could leverage CG to understand unseen medical images. To further validate this finding, we repeated the experiments on different MLLMs and obtained consistent results, confirming the universality of CG.
Building on these insights, we expanded the number of combinations and observed the changes in model generalization performance after deliberately disrupting CG, ultimately revealing that CG is a key factor driving the generalization of MLLMs. Furthermore, we explored the potential applications of CG and its performance across classification and detection tasks, finding that CG enhances MLLMs’ ability to handle medical scenarios with limited training data and improves their capacity for spatial awareness.
@@
Here are the key contributions of our work: 1) A VQA dataset, Med-MAT, has been constructed, providing a platform to explore the generalization of MLLMs on medical images. 2) Through this dataset, we observed that MLLMs in different architectures can utilize compositional generalization to understand unseen images and demonstrated that this is one of the main forms of generalization for medical MLLMs. 3) Finally, the real-world applicability of CG, along with its presence across detection and classification tasks, has been further explored, highlighting its potential to enhance data-efficient training and its broad applicability.
| Subset No. | 02 | 03 | 07 | 08 | 09 | 11 | 13 | 14 | 15 | 16 | 18 | 19 | 21 | 22 | 23 | 25 | 26 | 28 | 30 | 31 | 32 | 33 | 35 | 36 | 37 |
| Baseline | 21 | 47 | 40 | 25 | 26 | 27 | 28 | 24 | 22 | 24 | 25 | 23 | 49 | 26 | 25 | 24 | 49 | 30 | 49 | 21 | 49 | 20 | 25 | 23 | 19 |
@@ 2.1 Data Collection (Med-MAT)
Most existing datasets for MLLMs Zhang et al. (2023c); Li et al. (2024); Chen et al. (2024b), primarily VQA datasets, provide broad coverage but lack attribute annotations for individual samples, which are not suitable for CG exploration. To address this gap, we curated a large collection of image-text pairs to develop Med-MAT, ensuring that each sample is explicitly defined by MAT-Triplet.
Data Construction
@@
Med-MAT contains a total of 106 image-label pair medical datasets, sourced from various medical public challenges or high-quality annotated datasets. All datasets are categorized according to their MAT-Triplet, with data having identical elements grouped into a single subset (Figure 2). Labels are manually clustered to ensure that annotations with the same meaning are not repeatedly used. In total, Med-MAT covers 11 medical modalities
, 14 anatomical areas
, and 13 medical tasks
, hoping that it can spread across various medical tasks like a mat. (Data lists are shown in Appendix B)
Data Distribution
@@
All subsets are divided into training and test sets following their original distributions or using a 9:1 ratio. To ensure a fair comparison, each training set is limited to 3,000 samples 111Most datasets contain around 3,000 samples., with label balance maintained as much as possible. Any subset that cannot meet this requirement is treated as an OOD (out-of-distribution) dataset. For the test sets, we strictly balance the number of samples per label to ensure that the accuracy metric reliably reflects model performance.
QA Pairs Construction
@@
To enable MLLMs to directly train and test on Med-MAT, all image-label paired data were converted into a visual question-answering (VQA) format (Figure 3). Specifically, each subset was manually assigned 6 instructions to guide the MLLM in answering the subset task. For convenience, all samples were converted into single-choice questions with up to four options, and the remaining distractor options were randomly drawn from other labels within the subset. To mitigate potential evaluation biases arising from varying option counts, the ImageWikiQA dataset Zhang et al. (2024b), a non-medical dataset consisting of single-answer, four-option questions, was incorporated during the training.
2.2 Observation
@@
Experiment Setup We chose LLaVA-v1.5-7B-Vicuna Liu et al. (2023) as the base model due to its transparent pretraining process and minimal use of medical data, reducing the risk of knowledge leakage. Leveraging MLLM’s flexibility, we enabled task switching and generalization by adjusting prompts, streamlining generalization studies. Each experiment ran for 5 epochs on 8 A800 GPUs with a batch size of 32 and a learning rate of 5e-6.
Analysis
To access the generalization of MLLMs, we trained the baseline on all ID datasets to simulate Multi-task Training and separately trained on individual ID datasets to establish the Single-task Training as the control group. We then evaluated the models on all datasets. The results in Table 1 and 2 confirm that Multi-task Training outperformed Single-task Training on specific tasks and improved OOD prediction, suggesting certain data combinations enhance classification and identifying valuable combinations for medical tasks warrants further research. This observation leads to a research question (RQ):
What drives the generalization observed in MLLMs during Multi-task Training?
To address it, we aim to explore the generalization mechanism of MLLMs from the perspective of compositional generalization (CG).
@@ 3 Proof of Concept on CG
This section will prove the existence of CG in MLLMs, offering preliminary insights to address the RQ and providing support for further analysis.
| Related Combination | Target Subset | Baseline | Baseline+ | Trained | CG Helps | ||||
25 | 25 | 27 | ✓ |
47 | 46 | 50 | ✓ |
33 | 50 | 57 | ✓ |
49 | 53 | 51 | ✗ |
49 | 53 | 72 | ✓ |
37 | 33 | 39 | ✓ |
37 | 33 | 43 | ✓ |
37 | 31 | 43 | ✓ |
37 | 37 | 43 | ✓ |
37 | 37 | 43 | ✓ |
37 | 37 | 41 | ✓ |
49 | 48 | 51 | ✓ |
49 | 48 | 52 | ✓ |
49 | 48 | 51 | ✓ |
47 | 46 | 72 | ✓ |
30 | 21 | 49 | ✓ |
30 | 21 | 46 | ✓ |
33 | 28 | 28 | ✗ |
49 | 49 | 91 | ✓ |
49 | 50 | 81 | ✓ |
25 | 51 | 74 | ✓ |
49 | 52 | 52 | ✗ |
33 | 50 | 60 | ✓ |
25 | 25 | 36 | ✓ |
47 | 50 | 81 | ✓ |
47 | 50 | 71 | ✓ |
30 | 32 | 28 | ✗ |
30 | 32 | 35 | ✓ |
30 | 32 | 41 | ✓ |
30 | 32 | 42 | ✓ |
25 | 29 | 33 | ✓ |
25 | 29 | 33 | ✓ |
40 | 33 | 63 | ✓ |
40 | 33 | 63 | ✓ |
48 | 50 | 52 | ✓ |
48 | 50 | 55 | ✓ |
33 | 36 | 42 | ✓ |
23 | 33 | 32 | ✗ |
49 | 50 | 50 | ✗ |
49 | 51 | 62 | ✓ |
49 | 51 | 52 | ✓ |
49 | 51 | 58 | ✓ |
23 | 24 | 27 | ✓ |
3.1 Experiment Setup
To explore the existence of CG from a finer perspective, this section focuses on CG with only two MAT-Triplet elements varying while the third remains constant. Additionally, we identified specific Modality-Area pairs
@@ 3.1 Experiment Setup
, such as dermoscopy paired consistently with skin, which were treated as a special category. These 4 different fixed formats were classified into distinct Direction Types.
@@
We adhered to the training setup described in Section 2.2 and evaluated the model’s performance on the Target data. Baseline refers to the model without any training, while Trained refers to the model trained solely on Related data. To ensure that our conclusions are not influenced by the amount of training data, we randomly sampled an equal number of data from the Unrelated subsets, and this configuration is referred to as Baseline+.
3.2 Results
@@
Results are shown in Table 3 and it can be observed that almost all CG combinations are able to generalize to downstream tasks, highlighting that MLLMs can leverage CG to generalize Target data across all Direction Types. Besides that, since this experiment focused solely on two-element tuples, we further investigated three-element tuples in Appendix A.4, where we also observed similarly strong generalizations when obtaining MAT-Triplet elements from three different datasets.
@@
Take-away 1: MLLMs can leverage CG to understand unseen medical images.
| Related Combination | Target Subset | Qwen | Llama | ||||
@@
Take-away 2: Generalization arises in medical datasets in which at least partial MAT elements pre-exist during training.
3.3 Extending CG to other Backbones
LLaVA was selected as the baseline because its training data and processes are publicly available, ensuring minimal exposure to medical images and preventing bias in the integration of medical image knowledge into the MLLM. To ensure that the results are not affected by the training data or the visual encoder of LLaVA, we randomly sampled two combinations from each Direction Type to investigate CG on Qwen2-VL-7B Wang et al. (2024a) and Llama3.2-11B-Vision Meta AI (2024).
Qwen2-VL undergoes additional training on proprietary data based on ViT and incorporates a strategy to adjust the number of vision tokens according to resolution. Llama3.2-Vision, on the other hand, pretrains its own vision encoder from scratch using proprietary data. Thus, both models serve as a means to assess whether MLLMs with different training data and vision encoders can still leverage CG to understand unseen images, ensuring that CG is not merely an artifact of LLaVA’s data fitting or specific to its vision encoder.
Table 4 presents the experimental results, showing that both selected backbones exhibit a certain degree of generalization across most tasks. This suggests that despite differences in pre-train data and vision encoders, different MLLMs can still leverage CG to understand unseen images.
Take-away 3: CG persists across different MLLM backbones.
4 Scaling Combination in CG
After confirming that CG is indeed a form of generalization in MLLMs, we expanded the number of participating combinations to explore the generalizability of CG and examine its relationship with the generalization exhibited by Multi-task Training to address the RQ.
4.1 Experiment Setup
Two sub-questions have been defined to verify the applicability of CG in multiple data combinations and examine its role in Multi-task Training.
(Q1) While previous experiment on CG indicated that Unrelated combinations provide no benefit to Target data, can generalization arise when training incorporates more Unrelated combinations, simulating a multi-task scenario?
(Q2) Previous studies suggest that Multi-task Training generally promotes better generalization than single-task training. If the CG conditions in Multi-task Training are deliberately disrupted, will the resulting generalization effect be affected?
Selection Strategy
To ensure a balanced evaluation of Related and Unrelated combinations, Subset 03 and Subset 28 were chosen as Target datasets because they exhibit the most balanced ratios of Related to Unrelated subsets (13:11 for Subset 03 and 11:13 for Subset 28), making them ideal for providing a diverse range of compositions in the scale-up experiments.